OpenStudy (anonymous):

angle U and angle V are complementar angles.If angles U is thre leaa than twice angle V. what are the measures of the angles.

OpenStudy (anonymous):

since U and V r complementary angles, U + V = 90deg Let angleV = x deg then angleU = 2x -3 so x + (2x -3) = 90 3x - 3 = 90 3x = 90+3 x = 93/3 x = 31 angleV = 31deg angleU = 2(31) -3 = 62-3 = 59deg